OpenStudy (anonymous):

log(base 1/8) 1/9=x slove for x find the exact and approximate solutions

OpenStudy (anonymous):

\[\frac{\ln(\frac{1}{9})}{\ln(\frac{1}{8})}\]

OpenStudy (anonymous):

or if you want to be fancier you could write \[\frac{2\ln(3)}{3\ln(2)}\]

OpenStudy (anonymous):

then what comes after that

OpenStudy (anonymous):

a calculator

OpenStudy (anonymous):

lol thanks got it

OpenStudy (anonymous):

yw